OpenStudy (anonymous):

How can x+8=-x+7 be set up as a system of equations A. Y=x+8 Y=-x+7 B. y=x+8 Y=-3x+7 C. Y=x+5 Y=-2x+11 D. Y=8 Y=-3x+7

Directrix (directrix):

This equation looks like what you get after substituting for y for the system of equations: y = x + 8 y = -x + 7

Directrix (directrix):

If x + 8 = y and y = -x + 7, then by substitution, x + 8 = -x + 7